>How close would pre-stellar nebula of radius 200 AU have to be in order to have the same angular diameter as the 6 degree diameter of the Andromeda Galaxy as seen from earth? Calculate the apparent magnitude of the central star if it had a luminosity 10 times that of the sun.


1
Expert's answer
2021-10-21T16:33:18-0400

1)

"D=\frac a{\sin\frac{6°}2}=3821~AU,"

2)

"\lg\frac L{L_0}=0.4(M_0-M),"

"M=2.5."